# Naturalistic Computational Cognitive Science Towards generalizable models and theories that capture the full range of natural behavior

### v2

@@ 3.1 Behavior can be different in naturalistic experiments
Example 2: memory performance differs between naturalistic search and explicit memorization. When participants are explicitly instructed to memorize objects in a 3D home environment, their subsequent recall accuracy was significantly lower than when they incidentally encounter the same objects during visual search tasks (Helbing et al.,, 2020). This effect also extends to spatial memory—participants were able to place objects closer to their original positions in the search condition. These findings demonstrate that tasks focused explicitly on memory may not capture all aspects of how memory operates during naturalistic behavior. Thus, computational models of memory may need to be revised to account for these differences in naturalistic settings.
@@
Example 3: mice can learn $1000\times$1000 × faster during natural behavior, compared to two-alternative forced choice (2AFC) tasks. Recent research has shown that learning rates can dramatically differ between simplified laboratory tasks and more naturalistic settings (Rosenberg et al.,, 2021). When mice are tested in traditional 2AFC tasks, they typically require 10,000 training trials over 3-6 weeks to reach asymptotic performance of only about 67% accuracy. However, when allowed to freely explore a labyrinth (similar to natural burrow systems (Small,, 1901)), mice demonstrate remarkably accelerated learning (Rosenberg et al.,, 2021). In this setting, mice learned to navigate a complex maze with 63 T-junctions in just a few hours, optimizing their path to a reward after only about 10 reward experiences. The learning rate in the naturalistic setting was approximately $1000\times$1000 × times faster than in 2AFC paradigms. This dramatic difference in learning speed demonstrates how traditional simplified experimental paradigms may fundamentally underestimate an organism’s learning capabilities, and suggests that learning mechanisms may operate differently when they are engaged through more naturalistic behavior.
@@ 3.2 Neural systems can operate differently under naturalistic conditions
Example 1: the role of superior colliculus in eye movements vs. integrating sensory inputs for body control. Cisek and Green, (2024) argues that as we increased naturalism from head-fixed to head-free and body-free settings for studying monkeys, we expanded our theory of superior colliculus from controlling saccadic eye movement to generally integrating multimodal cues to guide bodily action-selection (Figure 4 B). That is, as we increased the naturalism of the experimental conditions, we arrived at a more complete model of superior colliculus. This illustrates the more general point, that overly focusing on computation as an input-output mapping neglects the fundamental fact that natural intelligence evolved not for single responses, but for closed-loop control in an environment (Cisek,, 1999). This perspectives yields a different interpretation of the system’s representations and processes. More generally, many researchers argue that cognition cannot be understood completely outside its embodiment and environment (e.g., Newen et al.,, 2018).
@@
Example 2: visual processing differs during natural viewing compared to passive viewing paradigms. One early visual response in the brain (the P100/M100) occurs about 100ms after a person’s eyes fixate on a new location (Vinje and Gallant,, 2000). However, this finding comes primarily from experiments where participants passively view images while fixating. In contrast, natural vision involves frequent active eye movements to sample information throughout a scene. When Amme et al., (2024) examined brain responses during more natural active viewing, they found that the P100/M100 response actually begins when the eye movement (saccade) begins, not when it ends at a new fixation point. Furthermore, the neural patterns during active viewing were anti-correlated to those seen during passive viewing, suggesting fundamentally different processing mechanisms. These findings reveal that simplified experimental paradigms, while valuable, may not capture all aspects of how visual processing operates during natural behavior.
@@
Example 3: model-based systems are used for learning moral judgments, but model-free systems are used for learning to avoid harming others. Despite moral reasoning being strongly associated with model-based learning in the prefrontal cortex, neural evidence shows a surprising shift when people actually learn to avoid harming others (Lockwood et al.,, 2020). Moral reasoning commonly engages model-based systems in the lateral prefrontal cortex (LFPC) (Spitzer et al.,, 2007; Carlson and Crockett,, 2018) with LPFC disruption leading to reduced norm compliance and enforcement (Knoch et al.,, 2006; Ruff et al.,, 2013). However, when participants perform tasks requiring them to learn moral behavior (rather than just reason about it), researchers instead observed neural activity consistent with model-free learning. As in §3.1, this illustrates that increasing naturalism of experiments (learning to take moral actions rather than simply reasoning about them), neural systems can operate in unexpected ways.
Here, we highlight examples where the challenges posed by increasing naturalism can engage computational mechanisms differently—which can help to disentangle underlying mechanisms, and achieve more generalizable understanding.
Example 1: Different learning mechanisms only yield distinct predictions under working memory load. One challenge of overly-simplified tasks is potential aliasing of different solutions, where many different computational approaches yield the same behavior. Collins, (2024) presents one example where a behavior ostensibly implemented as a standard reinforcement learning algorithm may instead be produced from simpler value-free associative learning combined with working memory. [^1] Whether this alternative counts as an alternative implementation of RL is orthogonal to our argument; our point is that the models yield distinct predictions in only some experimental regimes. They note that “Disentangling multiple processes requires considering more complex tasks to elicit differentiable behavior.” The more complex tasks in question simply increase the set of stimuli within learning blocks to a handful of objects, rather than only two or three—i.e., expanding the range of set sizes more clearly disentangles the learning processes. Thus, by restricting to minimal paradigms, we enable a system to use many solutions and prevent ourselves from discriminating between them. By contrast, by imposing a broad range of evaluations on the system (cf. Nau et al.,, 2024)—especially naturalistic evaluations that increase demands along different axes of task difficulty—the increased constraints actually make it easier to map the model on to natural intelligence (more constraints on the mapping make it less ambiguous), as is highlighted in the contravariance principle of Cao and Yamins, 2024b .
Example 2: fear conditioning. Traditional fear conditioning research offers fundamental insights into how humans learn associations between stimuli and aversive outcomes (Sehlmeyer et al.,, 2009; Maren,, 2001; Britton et al.,, 2014). Through carefully controlled laboratory studies, we have developed a rich mechanistic understanding of basic fear learning processes. These paradigms have proven especially valuable for studying certain types of naturally occurring fear learning, such as immediate aversive responses to foods (Riley and Tuck,, 1985) or traumatic events (Engelhard et al.,, 2008). However, real-world fear learning often involves additional temporal dynamics that complement these well-studied processes (Figure 4 C). For example, a student may develop a fear of exams after receiving a single poor grade weeks after taking a test, without requiring repeated negative experiences (Mobbs et al.,, 2021). This naturalistic learning involves delayed and sparse feedback. Thus, it relies on a computational mechanism for credit assignment (Sutton and Barto,, 2018)—a mechanism for crediting choices (e.g. studying or sleep) as responsible for future events (e.g. a poor grade). Naturalistic settings enable studying important functions such as mechanisms for credit assignment. Increased ecological validity can thereby increase the applicability of our findings; indeed, the gap between lab-based fear conditioning and real-world fear has been a barrier to treating pathological fears (Beckers et al.,, 2013; Scheveneels et al.,, 2016; Krypotos et al.,, 2018).
Example 3: seemingly-similar tasks that engage brain mechanisms differently can be reconciled by naturalistic models. There has been a long debate about whether perceptual processes like “oddity” tasks (identifying an object that differs from the others in a set) recruit the Medial Temporal Lobe (MTL). Some researchers (e.g Murray et al.,, 2007) suggest that MTL is recruited to discriminate among “complex” stimulus sets that visual cortex cannot discriminate on its own (thus MTL lesions selectively impair performance on those stimuli); other researchers have found no such interaction and suggested it stems from methodological issues (Buffalo et al.,, 1998; Suzuki,, 2009). Recently, stimulus-computable models have shed new light on this debate. Specifically, Bonnen et al., (2021); Bonnen and Eldridge, (2023) used pretrained convolutional vision models to directly process all the raw stimuli used in prior experiments, and evaluated whether the “oddity” stimulus could be distinguished from the rest of the set in the representation space of these models. By doing so, Bonnen et al. showed that purportedly “complex” stimuli that could be discriminated by the pretrained vision network alone did not require the MTL, while the complex stimuli that could not be distinguished by the vision model were precisely the ones that required MTL. This shows that the way that prior works had operationalized “complexity” was underspecified; differences in operationalization that got abstracted away in conceptual descriptions of the experiments likely drove the seemingly-conflicting findings. This example illustrates the value of building models that directly perform the same naturalistic tasks as the subjects, and testing these models on benchmarks that incorporate stimuli from many experimental paradigms. This naturalistic approach can elucidate important differences among experimental paradigms, and clarify conceptual understanding.
@@ 4 The (surprising) benefits of learning with naturalistic data
Figure 5: Overview of the benefits of learning with naturalistic data. (Figures reproduced from Doshi and Konkle,, 2023; Manning et al.,, 2020; Team et al.,, 2023.)
@@
A standard lesson in science is that we must simplify an experimental setting to better arrive at a causal conclusion. However, we have reviewed examples illustrating that more is different (Anderson,, 1972)—some properties of intelligence may only emerge in more “complicated” naturalistic settings. Here, we detail examples where naturalistic data itself plays an important role in producing the empirical phenomena of intelligence.
@@
Many results we discuss are drawn from modern AI research. One common trend is a positive interaction between learning-based systems and naturalistic data—learning-based systems can accommodate naturalistic data, and reciprocally, learning from naturalistic data results in qualitatively different generalization than learning in simpler settings. While we do not yet know the full implications for natural intelligence, these results open interesting questions and challenge prior assumptions. In particular, cognitive science should consider the limitations placed on our models when we fail to consider naturalistic data. (While we draw case studies from deep learning, we believe that similar factors could likewise affect generalization under other learning paradigms, such as program synthesis, or even evolution—and thus should be of interest even to those who do not work with deep learning models.)
4.1 Learning enables models that can operate on natural data
One surprising benefit of learning is that it can enable a system to operate over natural data. Until AlexNet (Krizhevsky et al.,, 2012), the prevailing wisdom was that one had to design useful representations of natural data for models. However, AlexNet and subsequent work showed that a model could learn useful representations for naturalistic data simply through learning to make predictions about those data. Surprisingly, these models learn representations that have notable correspondence to the internal representations in visual cortex (Yamins et al.,, 2014; Khaligh-Razavi and Kriegeskorte,, 2014). Here, we review some of this work and how learning-based methods can enable more flexible cognitive models that operate on natural data.
Learning-based vision models can operate on natural data and predict neural data (surprisingly) well. A rich history of work reveals how neural networks can learn internal representations from natural images that strongly correspond to neural data. A seminal example comes from Olshausen and Field, (1996), who showed that V1-like Gabor filters could emerge from a convolutional network optimized to transmit input while minimizing overall activation (essentially, a sparse autoencoder). This breakthrough demonstrated that V1-like receptive fields could arise from optimization for natural environmental statistics. Subsequently, Lee et al., (2007) demonstrated that V2-like convolutional filters could emerge from unsupervised optimization of sparse deep belief networks. Later, Yamins and DiCarlo, (2016) showed that deep convolutional networks trained for object classification developed intermediate layers that matched neural representations in V4 and IT better than any previous models. Strikingly, task-optimized networks trained to self-organize represent images in a latent space whose geometry corresponds closely to brain topography (Doshi and Konkle,, 2023, Figure 5 A). Thus, learning-based approaches not only enable effective visual processing but also can provide surprisingly accurate predictions of brain organization and function.
Integrating learning-based models with program synthesis enables flexible Bayesian models of human concept learning. Recently, Ellis, (2024) showed how language models can be combined with program synthesis to create more generalizable Bayesian models of human reasoning. Language models enable using natural language as a flexible hypothesis space with a prior derived from human data. By translating natural language hypotheses to executable programs for likelihood computation, such models can capture human learning dynamics in online concept acquisition tasks while continuing to predict human behavior robustly in new domains. This overcomes traditional limitations of hand-designed symbolic hypothesis spaces and pure program synthesis approaches, which can fail to transfer due to incomplete symbolic vocabularies—a manifestation of the frame problem (Shanahan,, 2004) in hypothesis space design.Integrating learning-based components with program synthesis thus enables Bayesian models that can flexibly adapt their hypothesis space while maintaining interpretable symbolic computation—a goal of many cognitive modelers.
@@ 4.2 Learning from naturalistic data can improve generalization
Studying how systems generalize is fundamental to cognitive science and AI. When studying generalization in simplified settings, a key assumption is that the simplifications do not alter the generalization problem. Here, we review studies from AI and neuroscience showing that variability in experience (even aloong seemingly-orthogonal axes) can alter learning and generalization—thus implying that studying models within more restricted settings may mislead us about the naturalistic computational problem.
@@
Example 1: Compositional generalization of image classifiers vs. agents. Naturalistic tasks can fundamentally change what models learn and how they generalize. Hill et al., (2020) trained two models on a vision-language grounding task, and tested their compositional generalization to held-out instructions. Both models were trained on the same language examples, and tested on the same held-out examples. However, one model was trained as an agent interacting with a simulated environment, while the other was simply trained trained to classify screenshots from that environment based on language instructions. The authors found that the interacting agent exhibited perfect compositional generalization to novel language utterances, while the image classifier generalized substantially worse. The authors also explored other factors, including generalization benefits of more naturalistic 3D (rather than 2D) environments, or egocentric embodiments. Their results illustrate how richer naturalistic settings can enhance the generalizability of learned solution. Thus, if we are interested in understanding generalization, we may need to build models that learn from appropriately naturalistic data.
@@
Example 2: Syntactic generalization is enhanced by variability in other structures’ fillers Structural generalization is of deep interest in linguistics. Recently, Misra and Mahowald, (2024) performed controlled experiments in which challenging linguistic constructions are systematically held out from language model training data, and showed that models trained on naturalistic language generalize to held out constructions by composing pieces of simpler constructions. Critically, this generalization depended on sufficient variability in the semantic fillers observed in the structures seen in training—thus, incorporating naturalistic variation in model training data can impact our theoretical inferences about generalization.
@@
Example 3: Rats raised in enriched environments. Analogously, neuroscientists have found that animals raised in more complex environments can be more skilled than those raised in simpler environments. Rats raised in enriched environments — with social interaction and/or more space or toys — show greater exploration, and improved learning and memory on novel tasks (Simpson and Kelly,, 2011). Thus, for natural as well as artificial intelligence, greater naturalistic variation in experience can alter the system’s learning in ways that impact our experimental and theoretical inferences.
@@
Example 1: Computer Vision. One early success of deep learning came in computer vision. Soon after AlexNet (Krizhevsky et al.,, 2012) achieved strong results on the ImageNet dataset (Deng et al.,, 2009), researchers showed that AlexNet’s feature could be repurposed to novel tasks like scene recognition or medical diagnosis (Donahue et al.,, 2014; Sharif Razavian et al.,, 2014; Litjens et al.,, 2017). This result was striking because features trained via supervision on fixed object classes nevertheless enabled transfer to novel objects or even different types of tasks. This pattern has continuend, and more recently researchers showed that the representations learned by these models could transfer from Imagenet 3D household navigation and dexterous tasks, sometimes achieving better performance than using “ground-truth” state (Parisi et al.,, 2022; Yuan et al.,, 2022). Beyond their useful representations, vision models trained with naturalistic data also develop mechanisms they were not trained for. For example, scene-oriented CNNs develop mechanisms for object detection (Zhou et al.,, 2014) and vision transformers develop mechanisms for segmentation (Caron et al.,, 2021)—both without an explicit training signal.
Example 2: Reinforcement learning. Likewise in reinforcement learning, researchers have found that agents trained on many (billions) of tasks exhibit strong “out-of-distribution” generalization on unknown tasks (Team et al.,, 2021), learning novel tasks as efficiently as humans (Team et al.,, 2023). Some algorithms can even generalize to collaborating with humans without any human data (Strouse et al.,, 2021). Recently, reinforcement learning algorithms for learning a modern variant of the successor representation (Gershman,, 2018; Carvalho et al.,, 2024) have been shown to develop skills for exploration and behavior without explicit training signals (Liu et al.,, 2024).
Example 3: Large language models. Perhaps the most striking example of this general phenomenon comes from large language models. These models are only trained to predict the next “token” (or word) to appear in a sequence. However, the distribution of internet language effectively includes a broad mixture of many tasks (Radford et al.,, 2019). When trained on these vast naturalistic data, language models develop mechanisms for disparate tasks such as modular arithithmic, solving word problems, etc. (Wei et al.,, 2022)—and even for adapting to new tasks from examples (Brown et al.,, 2020). Moreover, the representations learned by these models can transfer to superficially dissimilar downstream tasks (Lu et al.,, 2022), including playing video games (Reid et al.,, 2022). These examples illustrate how learning from a broad naturalistic distribution can induce many abilities—and can transfer to downstream tasks.

### v4

@@ 3.2 Behavior can be qualitatively different in naturalistic experiments
Example 1: memory performance differs between naturalistic search and explicit memorization. When participants are explicitly instructed to memorize objects in a 3D home environment, their subsequent recall accuracy was significantly lower than when they incidentally encountered the same objects during visual search tasks (Helbing et al., 2020). This effect also extends to spatial memory—participants were able to place objects closer to their original positions in the search condition. These findings demonstrate that tasks focused explicitly on memory may not capture all aspects of how memory operates during naturalistic behavior. Thus, computational models of memory may need to be revised to account for these differences in naturalistic settings.
@@
Example 2: mice can learn $1000\times$ faster during natural behavior, compared to two-alternative forced choice (2AFC) tasks. Recent research has shown that learning efficiency can dramatically differ between simplified laboratory tasks and more naturalistic settings (Rosenberg et al., 2021). When mice are tested in traditional 2AFC tasks, they typically require 10,000 training trials over 3-6 weeks to reach asymptotic performance of only about 67% accuracy. However, when allowed to freely explore a labyrinth (similar to natural burrow systems (Small, 1901)), mice demonstrate remarkably accelerated learning (Rosenberg et al., 2021); they learned to navigate a complex maze with 63 T-junctions in just a few hours, optimizing their path to a reward after only about 10 reward experiences. The time it took them to learn in the naturalistic setting was approximately $1000\times$ faster than in 2AFC paradigms. While it may not be meaningful to directly compare “learning rate” as a parameter between the settings, this clear and dramatic difference in learning efficiency demonstrates how traditional simplified experimental paradigms may fundamentally underestimate an organism’s learning capabilities, and suggests that learning mechanisms may operate differently when they are engaged through more naturalistic behavior.
Example 3: moral behavior can change when moving from textual vignettes to virtual reality. Recent research shows that behavioral choices can change when moving from text to virtual reality, by driving a shift from moral judgment (i.e. deciding whether things are moral) to moral action (Francis et al., 2017, see Figure 4 A). When participants were presented the classic trolley dilemma in text form, only 40% said they would pull a switch to divert the trolley, thus killing one person to save five others. However, when the same scenario was presented in virtual reality, where participants must simulate the action, 55% chose to pull the switch. The difference was even more pronounced in the “footbridge” variant of the dilemma. In the text version, only 10% of participants said they would push someone off a bridge to save five others. However, in virtual reality, 65% of participants choose to push the person. These findings demonstrate that people’s stated moral preferences in hypothetical scenarios can differ substantially from their actions in more naturalistic settings.
@@ 3.3 Neural systems can operate differently under naturalistic conditions
Example 1: the role of superior colliculus in eye movements vs. integrating sensory inputs for body control. Cisek and Green (2024) argues that as we increased naturalism from head-fixed to head-free and body-free settings for studying monkeys, we expanded our theory of superior colliculus from controlling saccadic eye movement to generally integrating multimodal cues to guide bodily action-selection (Figure 4 B). That is, as we increased the naturalism of the experimental conditions, we arrived at a more complete model of superior colliculus. This illustrates the more general point, that overly focusing on computation as an input-output mapping neglects the fundamental fact that natural intelligence evolved not for single responses, but for closed-loop control in an environment (Cisek, 1999). This perspectives yields a different interpretation of the system’s representations and processes. More generally, many researchers argue that cognition cannot be understood completely outside its embodiment and environment (e.g., Newen et al., 2018).
@@
Example 2: visual processing differs during natural viewing compared to passive viewing paradigms. One early visual response in the brain (the P100/M100) occurs about 100ms after a person’s eyes fixate on a new location (Vinje and Gallant, 2000). However, this finding comes primarily from experiments where participants passively view images while fixating. In contrast, natural vision involves frequent active eye movements to sample information throughout a scene. When Amme et al. (2024) examined brain responses during more natural active viewing, they found that the P100/M100 response actually begins when the eye movement (saccade) begins, not when it ends at a new fixation point. Furthermore, the neural patterns during active viewing were anti-correlated to those seen during passive viewing, suggesting fundamentally different processing mechanisms. These findings reveal that simplified experimental paradigms, while valuable, may not capture all aspects of how visual processing operates during natural behavior.
@@
Example 3: model-based systems are used for learning moral judgments, but model-free systems are used for learning to avoid harming others. Despite moral reasoning being strongly associated with model-based learning in the prefrontal cortex, neural evidence shows a surprising shift when people actually learn to avoid harming others (Lockwood et al., 2020). Moral reasoning commonly engages model-based systems in the lateral prefrontal cortex (LPFC) (Spitzer et al., 2007; Carlson and Crockett, 2018) with LPFC disruption leading to reduced norm compliance and enforcement (Knoch et al., 2006; Ruff et al., 2013). However, when participants perform tasks requiring them to learn moral behavior (rather than just reason about it), researchers instead observed neural activity consistent with model-free learning. As in §3.2, this illustrates that when we increase the naturalism of experiments (learning to take moral actions rather than simply reasoning about them), neural systems can operate in unexpected ways.
@@ 4 The (surprising) benefits of learning with naturalistic data
Figure 5: Overview of the benefits of learning with naturalistic data. (Figures reproduced from Doshi and Konkle, 2023; Manning et al., 2020; Adaptive Agent Team et al., 2023.)
@@
A standard lesson in science is that we must simplify an experimental setting to better arrive at a causal conclusion (e.g. Rust and Movshon, 2005). However, we have reviewed examples illustrating that more is different (Anderson, 1972)—some properties of intelligence may only emerge in more “complicated” naturalistic settings. Here, we detail examples where naturalistic data itself plays an important role in producing the empirical phenomena of intelligence.
@@
Many results we discuss are drawn from modern AI research. One common trend is a positive interaction between learning-based systems and naturalistic data—learning-based systems can accommodate naturalistic data, and reciprocally, learning from naturalistic data results in qualitatively different generalization than learning in simpler settings. While we do not yet know the full implications for natural intelligence, these results open interesting questions and challenge prior assumptions. In particular, cognitive science should consider the limitations placed on our models—and correspondingly, on the theories we instantiate in those models—when we fail to consider naturalistic data.
@@ 4.1 Learning from naturalistic data can improve generalization
Studying how systems generalize is fundamental to building theories in cognitive science and AI. In order to study generalization in simplified settings, we need to make the key assumption that the simplifications do not alter the generalization problem. Here, we review studies from AI and neuroscience showing that variability in experience (even along seemingly-orthogonal axes) can alter learning and generalization—thus implying that studying models within more restricted settings may mislead us about the naturalistic computational problem.
@@
Example 1: Compositional generalization of image classifiers vs. agents. Naturalistic tasks can fundamentally change what models learn and how they generalize. Hill et al. (2020) trained two models on a vision-language grounding task, and tested their compositional generalization to held-out instructions. Both models were trained on the same language set, and tested on the same held-out examples. However, one model was trained as an agent interacting with a simulated environment, while the other was simply trained to classify screenshots from that environment based on language instructions. Surprisingly, the interacting agent exhibited perfect compositional generalization to novel language utterances, while the image classifier generalized substantially worse. The authors also found generalization benefits of other naturalistic factors, including 3D (rather than 2D) environments, or egocentric embodiments. These results illustrate how richer naturalistic settings can enhance the generalizability of learned solutions. Thus, if we are interested in understanding generalization, we may need to build models that learn from appropriately naturalistic data.
@@
Example 2: Syntactic generalization is enhanced by variability in other structures’ fillers Structural generalization is of deep interest in linguistics. Recently, Misra and Mahowald (2024) performed controlled experiments in which challenging linguistic constructions are systematically held out from language model training data, and showed that models trained on naturalistic language generalize to held out constructions by composing pieces of simpler constructions. Critically, this generalization depended on sufficient variability in the semantic fillers observed in the structures seen in training—thus, incorporating naturalistic variation in model training data can impact our theoretical inferences about generalization.
@@
Example 3: Rats raised in enriched environments. Analogously, neuroscientists have found that animals raised in more complex environments can be more skilled than those raised in simpler environments. Rats raised in enriched environments — with social interaction and/or more space or toys — show greater exploration, and improved learning and memory on novel tasks (Simpson and Kelly, 2011). Thus, for natural as well as artificial intelligence, greater naturalistic variation in experience can alter the system’s learning in ways that impact our experimental and theoretical inferences.
@@
Example 1: Computer Vision. One early success of deep learning came in computer vision. Soon after AlexNet (Krizhevsky et al., 2012) achieved strong results on the ImageNet dataset (Deng et al., 2009), researchers showed that AlexNet’s features could be repurposed to novel tasks like scene recognition or medical diagnosis (Donahue et al., 2014; Sharif Razavian et al., 2014; Litjens et al., 2017). This result was striking because features trained via supervision on fixed object classes nevertheless enabled transfer to novel objects or even different types of tasks. This pattern has continued, and more recently researchers showed that the representations learned by these models could transfer from ImageNet to 3D household navigation and dexterous tasks, sometimes achieving better performance than using “ground-truth” state (Parisi et al., 2022; Yuan et al., 2022). Beyond their useful representations, vision models trained with naturalistic data also develop mechanisms they were not explicitly trained for. For example, scene-oriented CNNs develop mechanisms for object detection (Zhou et al., 2014) and vision transformers develop mechanisms for segmentation (Caron et al., 2021)—both without an explicit training signal.
Example 2: Reinforcement learning. Likewise in reinforcement learning, researchers have found that agents trained on many (billions) of tasks exhibit strong “out-of-distribution” generalization on unknown tasks (Open Ended Learning Team et al., 2021), learning novel tasks as efficiently as humans (Adaptive Agent Team et al., 2023). Some algorithms can even generalize to collaborating with humans without any human data (Strouse et al., 2021). Recently, reinforcement learning algorithms for learning a modern variant of the successor representation (Gershman, 2018; Carvalho et al., 2024) have been shown to develop skills for exploration and behavior without explicit training signals (Liu et al., 2024).
Example 3: Large language models. Perhaps the most striking example of this general phenomenon comes from large language models. These models are only trained to predict the next “token” (or word) to appear in a sequence. However, the distribution of internet language effectively includes a broad mixture of many tasks (Radford et al., 2019). When trained on these vast naturalistic data, language models develop mechanisms for disparate tasks such as modular arithmetic, solving word problems, etc. (Wei et al., 2022)—and even for adapting to new tasks from examples (Brown et al., 2020). Moreover, the representations learned by these models can transfer to superficially dissimilar downstream tasks (Lu et al., 2022), including playing video games (Reid et al., 2022). These examples illustrate how learning from a broad naturalistic distribution can induce many abilities—and can transfer to downstream tasks.